An ion with mass no. 56 contain 3 positive charge . If ion contain 30.4% more neutrons than the electrons??

Grade:12th pass

1 Answers

Ishika dammani
6 years ago
Let the number of electrons in ion = XNumber of proton or atomic number of element = X+3Number of neutrons in ion = X + 30.4% of X = X + 0.304 X=1.304 X Atomic mass of an element is sum of the number of neutrons and number of protons of the element.thus56 = (X+3) +(1.304 X) orX = 23Hence Atomic number of the element is 23+3 = 26, and the element is Fe,The ion is Fe3+